### v1

# Weaker Assumptions for Asymmetric Trust

Abstract
In distributed systems with asymmetric trust, each participant is free to make its own trust assumptions about others, captured by an asymmetric quorum system. This contrasts with ordinary, symmetric quorum systems and threshold models, where trust assumptions are uniformly shared among participants. Fundamental problems like reliable broadcast and consensus are unsolvable in the asymmetric model if quorum systems satisfy only the classical properties of consistency and availability. Existing approaches overcome this by introducing stronger assumptions. We show that some of these assumptions are overly restrictive, so much so that they effectively eliminate the benefits of asymmetric trust. To address this, we propose a new approach to characterize asymmetric problems and, building upon it, present algorithms for reliable broadcast and consensus that require weaker assumptions than previous solutions. Our methods are general and can be extended to other core problems in systems with asymmetric trust.
1 Introduction
Asymmetric trust models, such as those proposed by Damgård et al. [12], Alpos et al. [2], and Losa et al. [19] allow the development of distributed protocols in which each participant can operate under its own trust settings. These algorithms are built on top of asymmetric quorum systems, where each process independently defines its own quorums. The traditional consistency and availability properties must be satisfied by these systems, but as was shown by Li et al. [17], they are not enough to solve fundamental problems like reliable broadcast and consensus. Therefore, the models introduce additional assumptions on the quorums. The protocols for reliable broadcast and consensus of Alpos et al. [2] require the existence of a guild, a set of processes that contains a quorum for each of its members. Li et al. [17] and Losa et al. [19] identify similar conditions. These assumptions can be very strong and restrictive, and while they are sufficient to provide solutions to the aforementioned problems, we will show that they are not necessary. This leads to asymmetric algorithms that work under weaker assumptions, making such systems more flexible and usable.
We start by showing that current assumptions for algorithms such as reliable broadcast or consensus are so restrictive that it becomes unnecessary to use asymmetric trust models altogether. That is, given an asymmetric trust assumption that satisfies such requirements, we can derive an equivalent symmetric trust assumption such that no process will be worse off by adopting it. This result is surprising, considering that Alpos et al. [2] show that asymmetric trust is more expressive than symmetric trust.
To solve this, we introduce the notion of depth of a process and propose a new approach to characterize asymmetric problems based on it. We use this concept to provide more general definitions of unauthenticated, asymmetric, asynchronous, Byzantine reliable broadcast and binary consensus. We also propose new algorithms for these problems that forego the guild requirement in favor of weaker assumptions on depth.
The remainder of this document is organized as follows. Section 2 discusses related work in the asymmetric trust setting. Section 3 introduces the model and reviews key concepts that serve as the foundation for our work. We show in Section 4 that the current assumptions are so strong that they invalidate the benefits of asymmetric trust. Section 5 introduces the concept of depth. Sections 6 and 7 present algorithms for reliable broadcast and consensus that forego the guild assumption in favor of weaker depth assumptions. Finally, conclusions are drawn in Section 8.
2 Related Work
Threshold and symmetric quorum systems are fundamental tools in distributed computing, commonly used to establish the safety and liveness guarantees of various protocols [22, 20, 5, 10]. They constitute a crucial component of many systems in cloud computing [16, 15] and cryptocurrencies [13, 8, 9]. One drawback is that all participants must use the same quorum system, which restricts their freedom to choose who to trust. The asymmetric model, introduced by Damgård et al. [12] and further developed by Alpos et al. [2], represents a novel approach that generalizes the symmetric paradigm. It allows participants to make trust choices of their own, reflecting social connections or other information available to them outside the protocol. Each participant can specify its own quorum system, giving it the freedom to have its own trust assumptions. Amores-Sesar et al. [4] show that algorithms relying on symmetric quorums cannot be applied directly to the asymmetric model, as certain properties do not hold if the underlying quorums are asymmetric. This drives the need to adapt existing symmetric protocols and redefine protocol properties to fit the new model.
Existing work [2] introduces asymmetric versions of relevant primitives like reliable broadcast, binary consensus and common coin. Amores-Sesar et al. [4] extend this work by developing a DAG-based consensus protocol that relies on such primitives. Losa et al. [19] present another approach to model asymmetric trust. They forego the notion of fail-prone sets and strengthen the definition of quorums, by requiring each quorum to contain a quorum for each member. Li et al. [17] expand on this model and address the topic of which quorum properties are necessary or sufficient to solve core asymmetric problems such as consensus. Sheff et al. [25] propose a variant of the Paxos protocol that incorporates heterogeneous trust assumptions. Senn and Cachin [24] explore asymmetric trust in the crash-fault scenario and show that, in this setting, asymmetric trust reduces to symmetric trust.
Asymmetric trust introduces novel challenges to the design of distributed protocols. One significant hurdle is that of processes that have wrong friends, that is, the existence of processes that are correct and behave according to the underlying algorithm, but whose trust assumptions do not capture the actual faulty processes. Alpos et al. [2] use the term naive to refer to such participants and point out that their presence can lead to safety and liveness issues in asymmetric protocols. To address this, they introduce the concept of a guild, a set of processes that contains a quorum for each member, and provide algorithms for reliable broadcast and consensus that work correctly only in executions that have this structure. They point out that having a quorum system that only satisfies the consistency and availability properties can lead to liveness issues for some processes. In a slightly different model, Li et al. [17] prove that the traditional properties of quorum consistency and availability are not sufficient to solve reliable broadcast and consensus. They introduce a new property, called strong availability, very similar to the notion of a guild of Alpos et al. [2], and prove that this property is sufficient to solve the aforementioned problems. They do not comment on the necessity of such a property.
Asymmetric consensus protocols have also transcended into the blockchain space with XRP Ledger (https://xprl.org[11] and Stellar (https://stellar.org[21, 18] as main representatives. In the XRP Ledger, each participant declares its own assumptions by listing other participating nodes from which it considers votes [11, 3]. Stellar uses a similar approach, in which each participant keeps a list of participants it considers important and waits for a suitable majority of them to agree on a transaction before considering it settled [21, 19, 18].
3 Preliminaries
3.1 Model
All algorithms we propose are for the asynchronous unauthenticated setting. We will consider a system of $n$ processes $\mathcal{P}=\{p_{1},\dots,p_{n}\}$ that interact with each other by exchanging messages. A protocol for $\mathcal{P}$ consists of a collection of programs with instructions for all processes. They are presented using the event-based notation of Cachin et al. [10]. An execution starts with all processes in a special initial state; subsequently the processes repeatedly trigger events, react to events, and change their state through computation steps. A process that follows its protocol during an execution is called correct. A faulty process, also called Byzantine, may crash or deviate arbitrarily from its specification. We will assume that there is a low-level functionality for sending messages over point-to-point links between each pair of processes. This functionality is accessed through the events of sending and receiving a message. Point-to-point messages are authenticated and delivered reliably among correct processes. Here and from now on, the notation $\mathcal{A}^{*}$ for a system $\mathcal{A}\subseteq 2^{\mathcal{P}}$ , denotes the collection of all subsets of the sets in $\mathcal{A}$, that is, $\mathcal{A}^{*}=\{A^{\prime}|A^{\prime}\subseteq A,A\in\mathcal{A}\}$.
3.2 Asymmetric Trust Overview
We consider and expand the asymmetric trust model proposed by Alpos et al. [2]. We summarize here the necessary context for understanding this work; we refer the reader to their paper for a full presentation. In protocols with asymmetric trust, each participant is free to make its own individual trust assumptions about others, captured by an asymmetric quorum system. This contrasts with ordinary, symmetric and threshold quorum systems, where all participants share the same trust assumptions. Given a set of processes $\mathcal{P}$, an asymmetric fail-prone system $\mathbb{F}=[\mathcal{F}_{1},\dots,\mathcal{F}_{n}]$, where $\mathcal{F}_{i}$ represents the failure assumptions of process $p_{i}$, captures the heterogeneous model. Each $\mathcal{F}_{i}$ is a collection of subsets of $\mathcal{P}$ such that some $F\in\mathcal{F}_{i}$ with $F\subseteq\mathcal{P}$ is called a fail-prone set for $p_{i}$ and contains all processes that, according to $p_{i}$, may at most fail together in some execution [12]. We can, in turn, proceed to define asymmetric Byzantine quorum systems, denoted by $\mathbb{Q}$.
Definition 3.1.
An asymmetric Byzantine quorum system $\mathbb{Q}$ for $\mathds{F}$ is an array of collections of sets $\mathbb{Q}=[\mathcal{Q}_{1},\cdots,\mathcal{Q}_{n}]$ where $\mathcal{Q}_{i}\subseteq 2^{\mathcal{P}}$ for $i\in[1,n]$. The set $\mathcal{Q}_{i}\subseteq 2^{\mathcal{P}}$ is a symmetric quorum system of $p_{i}$ and any set $Q_{i}\in\mathcal{Q}_{i}$ is called a quorum for $p_{i}$. The system $\mathbb{Q}$ must satisfy the following two properties.
Consistency:
The intersection of two quorums for any two processes contains at least one process for which either process assumes that it is not faulty, i.e.,
$\forall i,j\in[1,n],\forall Q_{i}\in\mathcal{Q}_{i},\forall Q_{j}\in\mathcal{Q}_{j},\forall F_{ij}\in\mathcal{F}_{i}^{*}\cap\mathcal{F}_{j}^{*}:Q_{i}\cap Q_{j}\nsubseteq F_{ij}.$
Availability:
For any process $p_{i}$ and any set of processes that may fail together according to $p_{i}$, there exists a disjoint quorum for $p_{i}$ in $\mathcal{Q}_{i}$, i.e.,
$\forall i\in[1,n],\forall F_{i}\in\mathcal{F}_{i}:\exists Q_{i}\in\mathcal{Q}_{i}:F_{i}\cap Q_{i}=\emptyset.$
Given an asymmetric quorum system $\mathbb{Q}$ for $\mathbb{F}$, an asymmetric kernel system for $\mathbb{Q}$ is defined analogously as the array $\mathbb{K}=[\mathcal{K}_{1},\dots,\mathcal{K}_{n}]$ that consists of the kernel systems for all processes in $\mathcal{P}$. A set $K_{i}\in\mathcal{K}_{i}$ is called a kernel for $p_{i}$ and for each $K_{i}$ it holds that $\forall Q_{i}\in\mathcal{Q}_{i},\ K_{i}\cap Q_{i}\neq\emptyset$, that is, a kernel intersects all quorums of a process.
Given an asymmetric fail-prone system $\mathbb{F}$, there will exist a valid asymmetric quorum system for $\mathbb{F}$ if and only if $\mathbb{F}$ satisfies the $B^{3}$ condition [12, 2]. This property is defined as follows.
Definition 3.2 ($B^{3}$-condition).
An asymmetric fail-prone system $\mathbb{F}$ satisfies the $B^{3}$-condition, abbreviated as $B^{3}(\mathbb{F})$, whenever it holds that
$\forall i,j\in[1,n],\forall F_{i}\in\mathcal{F}_{i},\forall F_{j}\in\mathcal{F}_{j},\forall F_{ij}\in{\mathcal{F}_{i}}^{*}\cap{\mathcal{F}_{j}}^{*}:\,\mathcal{P}\not\subseteq F_{i}\cup F_{j}\cup F_{ij}$
If $B^{3}(\mathbb{F})$ holds, then the canonical quorum system, defined as the complement of the asymmetric fail-prone system, is a valid asymmetric quorum system. This is similar to what happens in the symmetric case, where if the fail-prone system $\mathcal{F}$ satisfies the $Q^{3}$ property, the existence of a valid symmetric quorum system $\mathcal{Q}$ is guaranteed [14].
During an execution, the set of processes that fail is denoted by $F$. The members of $F$ are unknown to the processes and can only be identified by an outside observer and the adversary. A process $p_{i}$ correctly foresees $F$ if $F\in\mathcal{F}_{i}^{*}$, that is, $F$ is contained in one of its fail-prone sets. Based on this information, it is possible to classify processes in three categories.
Faulty:
a faulty process, i.e., $p_{i}\in F$;
Naive:
a correct process $p_{i}$, i.e. $p_{i}\notin F$, where $F\notin\mathcal{F}^{*}_{i}$; or
Wise:
a correct process $p_{i}$, i.e. $p_{i}\notin F$, where $F\in\mathcal{F}^{*}_{i}$
Alpos et al. [2] show that naive processes may disrupt the safety and liveness guarantees of some protocols. In order to formalize this notion, they introduced the concept of a guild which is central for many of the algorithms they propose (e.g., reliable broadcast, binary consensus).
Definition 3.3.
A guild is a set of wise processes that contains one quorum for each of its members. Formally, a guild $\mathcal{G}$ for $\mathds{F}$ and $\mathds{Q}$, for an execution with faulty processes $F$, is a set of processes that satisfies the following two properties.
Wisdom:
$\mathcal{G}$ is a set of wise processes, that is,
$\forall p_{i}\in\mathcal{G}:F\in\mathcal{F}_{i}^{*}.$
Closure:
$\mathcal{G}$ contains a quorum for each of its members, that is,
$\forall p_{i}\in\mathcal{G}:\exists Q_{i}\in\mathcal{Q}_{i}:Q_{i}\subseteq\mathcal{G}.$
Alpos et al. [2] provide reliable broadcast and binary consensus implementations which function only in executions that contain a guild. In relation to this, Definition 3.4 presents the notion of a tolerated system, originally introduced by Alpos et al. [2]. It is built on top of tolerated sets, which are sets that contain all elements outside of a guild. This portrays the resilience of the trust assumptions, as even when all processes in one of the tolerated sets fail, there is still a guild in the system. This structure is used by Alpos et al. [2] and Zanolini [26] in their construction of several asymmetric protocols.
Definition 3.4 (Tolerated system [2]).
Given an asymmetric Byzantine quorum system $\mathbb{Q}$ and an execution with faulty processes $F$, a set of processes $T$ is called tolerated (by $\mathbb{Q}$) if a non-empty guild $\mathcal{G}$ for $F$ and $\mathbb{Q}$ exists such that $T=\mathcal{P}\setminus\mathcal{G}$.
The tolerated system $\mathcal{T}$ of an asymmetric Byzantine quorum system $\mathbb{Q}$ is the maximal collection of tolerated sets, where $F$ ranges over all possible executions.
Lemma 3.1, which is crucial for the construction presented in Section 4, shows that if there is an asymmetric fail-prone system $\mathbb{F}$ such that $B^{3}(\mathbb{F})$ holds, then $Q^{3}(\mathcal{T})$ also holds. This means that $\mathcal{T}$ is a valid symmetric fail-prone system [14].
Lemma 3.1 ([2]).
Let $\mathbb{Q}$ be an asymmetric Byzantine quorum system among processes $\mathcal{P}$ with asymmetric fail-prone system $\mathbb{F}=\overline{\mathbb{Q}}$, i.e., such that $\mathbb{Q}$ is a canonical asymmetric Byzantine quorum system, and let $\mathcal{T}$ be the tolerated system of $\mathbb{Q}$. If $B^{3}(\mathbb{F})$ holds, then $\mathcal{T}$ is a valid symmetric fail-prone system, i.e., $Q^{3}(\mathcal{T})$ holds.
3.3 Quorum-based Algorithms
In the following sections, we will give results regarding the existence of algorithms to solve problems such as reliable broadcast and consensus. To do this, we must first define what we mean by an algorithm. We will consider all algorithms to be quorum-based algorithms. Definition 3.5 formalizes this notion. It is derived from the definitions used by Losa et al. [19] and Li et al. [17] , but it is adapted to fit the asymmetric model of Alpos et al. [2]. We use the term initiator of the execution to refer to the sole process that triggers the execution of an algorithm. In our work, this only applies within the context of reliable broadcast, where the initiator is the process that broadcasts its value. In consensus, there is no single initiator because all processes must propose a value.
Definition 3.5 (Quorum-based algorithm).
Given a distributed algorithm $A$, processes $\mathcal{P}$, and asymmetric quorum system $\mathbb{Q}$, we will say that $A$ is quorum-based if the following properties hold
  • In any execution where a correct process $p_{i}$ issues a response, there is a quorum $Q_{i}\in\mathcal{Q}_{i}$ such that $p_{i}$ received a message from each member of $Q_{i}$.
  • If the state of a correct process $p_{i}$ changes upon receiving a message from some process $p_{j}$, then there is a $Q_{i}\in\mathcal{Q}_{i}$ such that $p_{j}\in Q_{i}$ or $p_{j}$ is the initiator of the execution of $A$.
  • For every execution where a correct process $p_{i}$ issues a response, there exists an execution where $p_{i}$ only receives messages from the members of one of its quorums $Q_{i}\in\mathcal{Q}_{i}$ and the initiator and $p_{i}$ issues the same response.
4 Guild Assumptions are not Asymmetric
In the asymmetric trust setting, Li et al. [17] show that a quorum system satisfying the classical properties of consistency and availability alone does not suffice to solve fundamental problems such as reliable broadcast or consensus. The heterogeneous views of the system make it necessary to assume more structure. Different approaches have reached similar conclusions about the assumptions needed to address these problems under asymmetric trust. Alpos et al. [2] require the existence of a guild, while Li et al. [17] propose that the system must satisfy the strong availability property. This is closely related to the idea of a guild, as it also requires the existence of a well-behaved group of processes that includes a quorum for every participant. In this section, we examine the implications of these assumptions,focusing on the model proposed by Alpos et al. [2].
Guilds are a strong assumption, essentially requiring all members to have common beliefs, which goes against the idea of asymmetric trust. We show that for any asymmetric Byzantine fail-prone system $\mathbb{F}$, if a guild exists in an execution, then we can build an equally valid symmetric trust assumption $\mathcal{F}$ from $\mathbb{F}$. This means that, in such cases, asymmetric trust reduces to symmetric trust.
In the crash-fault model, Senn and Cachin [24] already show that given an asymmetric fail-prone system $\mathbb{F}$, it is possible to construct a symmetric fail-prone system $\mathcal{F}$ from $\mathbb{F}$ such that if all processes adopt $\mathcal{F}$ as their trust assumption, no process will be worse off. This result implies that there is no need for asymmetric trust in that setting. We extend this by showing that a similar scenario occurs in the Byzantine fault model for all executions with a guild. This effectively invalidates the advantages of asymmetric trust for the algorithms proposed by Alpos et al. [2] which rely on the existence of a guild.
We now show how to construct a symmetric fail-prone system $\mathcal{F}$ from an asymmetric fail-prone system $\mathbb{F}$, such that in all executions with a guild, $\mathcal{F}$ is as expressive as $\mathbb{F}$. We will build it based on the notion of the tolerated system, presented in Section 2 and used by Alpos et al. [2] and Zanolini [26] in their consensus protocols.
Definition 4.1 (Tolerated Symmetric Fail-Prone System).
Let $\mathbb{F}$ be an asymmetric fail-prone system that satisfies $B^{3}(\mathbb{F})$, $\mathbb{Q}$ be its associated canonical quorum system and $\mathcal{T}$ be the tolerated system derived from $\mathbb{Q}$ according to Definition 3.4. We interpret $\mathcal{T}$ as a symmetric fail-prone set and denote it by $\hat{\mathcal{F}}$, that is, $\hat{\mathcal{F}}=\mathcal{T}$, with a one-to-one correspondence between the tolerated sets and fail-prone sets.
From Lemma 3.1 we know that the tolerated symmetric fail-prone system satisfies the $Q^{3}$ property and therefore there will be at least one valid symmetric Byzantine quorum system associated to it.
Lemma 4.1.
Let $\mathbb{F}$ be an asymmetric fail-prone system that satisfies $B^{3}(\mathbb{F})$, $\mathcal{T}$ be the tolerated system derived from $\mathbb{F}$ and $\hat{\mathcal{F}}$ be the tolerated symmetric fail-prone system derived according to Definition 4.1. Given an execution with a guild $\mathcal{G}$ and faulty processes $F$, then $F\in\hat{\mathcal{F}}$.
Proof.
By Definition 3.4, there exists a tolerated set $T=\mathcal{P}\setminus\mathcal{G}$ such that $T\in\mathcal{T}$. Since $F\cap\mathcal{G}=\emptyset$ we know that $F\subseteq T$ and consequently there exists a fail-prone set $F^{\prime}\in\hat{\mathcal{F}}$ such that $T=F^{\prime}$. Therefore, the set of faulty processes will be considered by the fail-prone system $\hat{\mathcal{F}}$. ∎
Theorem 4.1.
Let $\mathbb{F}$ be an asymmetric fail-prone system that satisfies $B^{3}(\mathbb{F})$. Given an execution with faulty processes $F$ and at least one guild $\mathcal{G}$, it is possible to derive a symmetric fail-prone system $\mathcal{F}$ from $\mathbb{F}$ such that $F\in\mathcal{F}^{*}$.
Proof.
[…]
Theorem 4.1 shows that for executions with a guild, every process is at least as well off using the symmetric quorum system $\mathcal{F}$ instead of $\mathbb{F}$. For a wise process $p_{i}$ the situation will remain the same regardless of the quorum system chosen. With the asymmetric system $F\in\mathcal{F}^{*}_{i}$ will hold, while with the symmetric alternative $F\in\mathcal{F}^{*}$. For naive processes the situation improves, as now the faulty processes will be considered in their trust assumption. Thus, there is no reason for a process not to adopt the derived symmetric fail-prone system $\mathcal{F}$. Building $\mathcal{F}$ can be computationally expensive because it requires computing $\mathcal{T}$. However, existing asymmetric consensus protocols [2, 26] also require computing $\mathcal{T}$, so our approach is no more costly than current solutions.
Hence, if a guild is needed to solve a problem in the asymmetric setting, there exists a way to circumvent this and use any existing symmetric algorithm while obtaining the same guarantees that an asymmetric algorithm would provide. Therefore, if a problem can only be solved if a guild exists, the interest in using asymmetric trust to solve it decreases. This motivates the search for better algorithms for problems such as reliable broadcast and consensus, where the only known asymmetric algorithms require a guild.
One limitation of this result is that it assumes knowledge of the trust assumptions for all processes ($\mathbb{F}$). Although this assumption holds in the Alpos et al. [2] model, it is not required in other asymmetric trust models [19]. We also note that knowledge of the faulty processes $F$ is not required to derive the symmetric fail-prone system presented in Theorem 4.1.
5 Depth
In the symmetric world, processes are classified as correct or faulty. This suffices to express the properties of algorithms and capture all potential states of a process. In the asymmetric world this is not enough, a process might be correct but have incorrect trust assumptions. Alpos et al. [2] address this problem by introducing the notions of naive, wise, and guilds.
Existing protocols for reliable broadcast and consensus in the Alpos et al. [2] model require a guild to ensure safety and liveness. The results of Section 4, coupled with the inherent rigidity of having this kind of assumptions, motivate the search for new asymmetric algorithms that forego such assumptions. In order to do this, we introduce a new way to describe the quality of processes, based on how much they can rely on other processes during an execution. We denote this concept as the depth of a process. We then show how it can be used to characterize asymmetric problems in a more fine-grained manner. Definition 5.1 introduces this notion. This approach allows to generalize the concept of wise process or guild member and introduces more expressivity to the system.
Definition 5.1 (Depth of a process).
For an execution, we recursively define the notion of a correct process having depth $d$ as follows
  • Any correct process $p_{i}$ has depth 0.
  • Additionally, a correct process $p_{i}$ has depth $d\geq 1$ if it has a quorum such that all processes contained in it have depth at least d1d-1, i.e.,
[…]
Note that a process with maximal depth $d$ also has depth $d^{\prime}$ for all $0\leq d^{\prime}\leq d$. We will focus on the maximal depth of a process. In the following sections, any reference to the depth of a process should be understood as its maximal depth. Following the terminology of Alpos et al. [2], naive processes have depth 0, wise processes have depth at least 1, and processes in a guild have infinite depth.
In Sections 6 and 7 we show how the properties of protocols can be adapted to take into account the depth of processes. We modify them to show that some property must hold for all processes with depth at least $d$. For example, the totality property of reliable broadcast is updated to state that all processes with depth $d$ eventually deliver a message.
6 Reliable Broadcast
Section 4 shows that if an execution includes a guild, we can convert asymmetric trust assumptions into symmetric ones. This means that if a problem can only be solved in executions with a guild, the interest in using asymmetric trust to solve it decreases. So far, the only known protocols for reliable broadcast and consensus in the Alpos et al. [2] model require the existence of a guild. This motivates the search for new algorithms that work without relying on guild assumptions for these problems. We start this search with the Byzantine reliable broadcast problem. Our approach weakens the trust assumptions needed and eliminates the need for a guild. Definition 6.1 presents a way to characterize reliable broadcast based on the depth of processes. Its properties are specific to processes that have at least a certain maximal depth.
Definition 6.1 (Depth-characterized asymmetric Byzantine reliable broadcast).
A protocol for asymmetric Byzantine reliable broadcast with sender $p_{s}$ and depth $d$, shortened as RB[$d$], defined through the events dar-broadcast(m) and dar-deliver(m) , satisfies the following properties:
  • Validity: If a correct process $p_{s}$ dar-broadcasts a message $m$, then all processes with depth $d$ eventually dar-deliver $m$.
  • Consistency: If some process with depth $d$ dar-delivers $m$ and another process with depth $d$ dar-delivers $m^{\prime}$, then $m=m^{\prime}$.
  • Integrity: Every process with depth $d$ dar-delivers $m$ at most once. Moreover, if the sender $p_{s}$ is correct and the receiver has depth $d$, then $m$ was previously dar-broadcast by $p_{s}$.
  • Totality: If a process with depth $d$ dar-delivers some message, then all processes with depth $d$ eventually dar-deliver a message.
The reliable broadcast proposed by Alpos et al. [2] is closely related to RB[$\infty$], although not as general. RB[$\infty$] guarantees a solution for all processes with depth $\infty$, whereas the protocol by Alpos et al. [2] only guarantees a solution for those processes with depth $\infty$ that also belong to the maximal guild. That is, processes with depth $\infty$ that are members of a guild, but not of the maximal guild, are not guaranteed to receive a solution.
A natural question arises, which is the minimum depth needed to implement the different primitives. A low depth not only corresponds to weaker asymmetric assumptions, but it also reflects, to a certain level, the communication complexity of an algorithm.
Lemmas 6.1 and 6.2 guide the search for an algorithm. Lemma 6.1 shows that if an algorithm solves RB[$s$] then it also solves RB[$s^{\prime}$] for all $s^{\prime}\geq s$. This simplifies the search for a solution by reducing it to finding an algorithm that works for the minimal value of $s$. Lemma 6.2 shows that there are no algorithms that can solve RB[1], therefore, we must search for a protocol that solves the problem for processes with a depth of at least 2.
Lemma 6.1.
If a quorum-based algorithm $A$ solves RB[d] then it solves RB[d] for all $\textit{d'}>\textit{d}$
Proof.
Suppose that an algorithm $A$ solves RB[d]. Denote by $S$ the set of all processes that have depth d and by $S^{\prime}$ the set of all processes that have depth d’. Since $S^{\prime}\subseteq S$ then all properties of RB[d] will be satisfied for the processes in $S^{\prime}$ after executing $A$. Therefore $A$ also solves RB[d].
Lemma 6.2.
There is no quorum-based algorithm that solves RB[1]
Proof.
Consider a system with processes $p_{1},p_{2},p_{3},p_{4},p_{5}$ and $p_{6}$ and the following fail-prone system.
$\mathbb{F}_{D}:\quad\begin{aligned} \mathcal{F}_{1}&=\{\{1,2,5,6\},\{1,2,3\}\}\\ \mathcal{F}_{2}&=\{\{1,2,5,6\},\{1,2,3\}\}\\ \mathcal{F}_{3}&=\{\{1,2,4\}\}\\ \mathcal{F}_{4}&=\{\{1,2,3\}\}\\ \mathcal{F}_{5}&=\{\{1,2,4\}\}\\ \mathcal{F}_{6}&=\{\{1,2,4\}\}\end{aligned}$
It satisfies the $B^{3}$ property and therefore the following canonical quorum system is valid.
$\mathbb{Q}_{D}:\quad\begin{aligned} \mathcal{Q}_{1}&=\{\{3,4\},\{4,5,6\}\}\\ \mathcal{Q}_{2}&=\{\{3,4\},\{4,5,6\}\}\\ \mathcal{Q}_{3}&=\{\{3,5,6\}\}\\ \mathcal{Q}_{4}&=\{\{4,5,6\}\}\\ \mathcal{Q}_{5}&=\{\{3,5,6\}\}\\ \mathcal{Q}_{6}&=\{\{3,5,6\}\}\end{aligned}$
Let the processes $p_{1}$, $p_{2}$, $p_{3}$, $p_{4}$ be correct and $p_{5}$, $p_{6}$ be Byzantine. Since $p_{1}$ and $p_{2}$ have a quorum composed only of correct processes ($\{3,4\}$) they have depth 1. We now show an execution in which $p_{1}$ dar-delivers while $p_{2}$ does not, which breaks the totality property of RB[1].
Firstly consider an execution $E_{0}$ in which $p_{5}$ dar-broadcasts value 0, all processes behave according to the algorithm. Processes $p_{1}$ and $p_{2}$ will deliver value 0. From Definition 3.5 we know that there is an equivalent execution $E$ where $p_{1}$ and $p_{2}$ deliver the same value and only receive messages from the processes in their quorum $\{4,5,6\}$ and $p_{5}$.
Now consider an execution $E^{\prime}$ in which processes $p_{5}$ and $p_{6}$ behave towards $p_{4}$ like in $E$. Note that $p_{4}$ will behave like in $E$ since at every step it will receive the same messages from the same processes as it did in $E$. Let $p_{5}$ and $p_{6}$ also behave towards $p_{2}$ like in $E$ and send no messages towards all other processes. Process $p_{3}$ does not begin the execution since it never receives any message from the initiator (Byzantine) nor any member of one of its quorums (processes 3, 5 or 6). Process $p_{2}$ will deliver value 0 as it receives the same messages from the same processes as in $E$. On the contrary, process $p_{1}$ will not deliver any value since it will never receive messages from all processes in one of its quorums, which would contradict Definition 3.5. This comes from the fact that $p_{3}$ will never send any message and that $p_{5}$ and $p_{6}$ are Byzantine and will not send it any message. Therefore, $p_{2}$ delivers a value but $p_{1}$ does not, which breaks the totality property of RB[1]. ∎
Algorithm 1 presents a solution for RB[3]. Every process waits to receive a quorum of Readyafterecho messages associated to the same message $m$ and round $r$ before delivering $m$. A process with depth 3 will only receive such a quorum if there exists a set of processes $Q$ (which form a quorum for a wise process) that can attest that the sender indeed sent the value $m$. Since there will be an attesting quorum for every process that delivers, and since quorums for wise processes intersect in at least one correct process, we can deduce that all processes with depth 3 that deliver a value $m$ will deliver the same value. In addition, we use a technique similar to Bracha’s kernel amplification [7] to ensure that if a valid process delivers then all valid processes will deliver. The arrays in lines 1,  1,  1, and 1 are hashmaps, so even though they are depicted as having infinite size they are actually sparsely populated. The algorithm has a latency of 3 asynchronous rounds in the good case and 5 in the bad case. Its correctness is proved in Lemma 6.3. The existence of an algorithm to solve RB[2] remains an open question.
State
$\textit{sentecho}\leftarrow\textsc{false}$: indicates whether $p_{i}$ has sent echo
$\textit{echos}\leftarrow[\bot]^{n}$: collects the received echo messages from other processes
$\textit{sentrae}\leftarrow[\textsc{false}]^{\infty}$: indicates whether $p_{i}$ has sent readyafterecho in round n, a hashmap
$\textit{sentrar}\leftarrow[\textsc{false}]^{\infty}$: indicates whether $p_{i}$ has sent readyafterready in round n, a hashmap
$\textit{readysafterecho}\leftarrow[\bot,\bot]^{\infty\times n}$: collects readyafterecho messages from other processes, a hashmap
$\textit{readysafterready}\leftarrow[\bot,\bot]^{\infty\times n}$: collects readyafterready messages from other processes, a hashmap
$\textit{delivered}\leftarrow\textsc{false}$: indicates whether $p_{i}$ has delivered a message
upon invocation $\textsl{dar-broadcast}(m)$ do
send message $\textsc{send},{m}]$ to all $p_{j}\in\mathcal{P}$
upon receiving a message $\textsc{send},{m}]$ from $p_{s}$ such that $\neg\textit{sentecho}$ do
$\textit{sentecho}\leftarrow\textsc{true}$
send message $\textsc{echo},{m}]$ to all $p_{j}\in\mathcal{P}$
upon receiving a message $\textsc{echo},{m}]$ from $p_{j}$ do
if $\textit{echos}[j]=\bot$ then
$\textit{echos}[j]\leftarrow m$
upon exists $m\not$$\bot$ such that $\{p_{j}\in\mathcal{P}|\textit{echos}[j]=m\}\in\mathcal{Q}_{i}$ and $\neg\textit{sentrae[1]}$ do // a quorum for $p_{i}$
$\textit{sentrae[1]}\leftarrow\textsc{true}$
send message $\textsc{readyafterecho, 1},{m}]$ to all $p_{j}\in\mathcal{P}$
upon exists $m\not=\bot,r>0$ such that $\{p_{j}\in\mathcal{P}|\textit{readysafterecho}[r][j]=m\}\in\mathcal{K}_{i}$ and $\neg\textit{sentrar[r]}$ do
$\textit{sentrar[n]}\leftarrow\textsc{true}$
send message $\textsc{readyafterready},{r,m}]$ to all $p_{j}\in\mathcal{P}$
upon receiving a message $\textsc{readyafterecho},{r,m}]$ from $p_{j}$ do
if $\textit{readysafterecho}[r][j]=\bot$ then
$\textit{readysafterecho}[r][j]\leftarrow m$
upon receiving a message $\textsc{readyafterready},{r,m}]$ from $p_{j}$ do
if $\textit{readysafterready}[r][j]=\bot$ then
$\textit{readysafterready}[r][j]\leftarrow m$
upon exists $m\not=\bot,r>1$ such that $\{p_{j}\in\mathcal{P}|\textit{readyafterready}[r][j]=m\}\in\mathcal{Q}_{i}$ and $\neg\textit{sentrae[r]}$ do
$\textit{sentrae[r+1]}\leftarrow\textsc{true}$
send message $\textsc{readyafterecho},{r+1,m}]$ to all $p_{j}\in\mathcal{P}$
upon exists $m\not=\bot,r>0$ such that $\{p_{j}\in\mathcal{P}|\textit{readysafterecho}[r][j]=m\}\in\mathcal{Q}_{i}$ and $\neg\textit{delivered}$ do
$\textit{delivered}\leftarrow\textsc{true}$
output dar-deliver(m)
Algorithm 1 Depth 3 asymmetric reliable broadcast with sender $p_{s}$ (RB[3]) (process $p_{i}$)
Lemma 6.3.
Algorithm 1 solves RB[3].
Proof.
Validity: This follows directly from the structure of the protocol. If a correct process broadcasts a value $m$ then all processes with depth 0 will send $\textsc{Echo},{m}]$ messages. Therefore all processes with depth 1 will receive such messages from at least one of their quorums. Then all processes with depth 1 will send $\textsc{Readyafterecho},{1,m}]$ messages. In the next round, all processes with depth 2 will receive such messages from one of their quorums and will deliver the value $m$. Since processes with depth 3 also have depth 2 this means that all processes with depth 3 will deliver $m$, satisfying the validity property.
Consistency: Suppose a process $p_{i}$ with depth 3 dar-delivered a value $m_{i}$ and another process with depth 3 $p_{j}$ dar-delivered a value $m_{j}$. By Lemma 6.6 this means that there exists a process with depth 1 $p_{i}^{\prime}$ such that it received $\textsc{Echo},{m_{i}}]$ messages from one of its quorums $Q_{i^{\prime}}\in\mathcal{Q}_{i^{\prime}}$ and another process with depth 1 $p_{j}^{\prime}$ that received $\textsc{Echo},{m_{j}}]$ messages from one of its quorums $Q_{j^{\prime}}\in\mathcal{Q}_{j^{\prime}}$. Since $Q_{i}^{\prime}\cap Q_{j}^{\prime}\nsubseteq F$ there is at least one correct process that sent $\textsc{Echo},{m_{i}}]$ and $\textsc{Echo},{m_{j}}]$ messages. But since a correct process will only send one type of echo messages, we conclude that $m_{i}=m_{j}$.
Integrity: The variable delivered guarantees that no correct process delivers multiple times. Moreover, if the sender $p_{s}$ is correct, the only message $m^{\prime}$ such that $p_{s}$ sends a message $\textsc{send},{m}]^{\prime}$ is $m$, thus $m$ was previously dar-broadcast by $p_{s}$.
Totality: Suppose that a process with depth 3 $p_{i}$ has dar-delivered a value $m$. Then it has obtained $\textsc{Readyafterecho},{n,m}]$ messages from one of its quorums $Q_{i}\in mathcal{Q}_{i}$. Consider any other process with depth 1 $p_{j}$. Since $p_{i}$ and $p_{j}$ are both wise, it holds $F\in\mathcal{F}_{i}^{*}$ and $F\in\mathcal{F}_{j}^{*}$, which implies $F\in\mathcal{F}_{i}^{*}\cap\mathcal{F}_{j}^{*}$. Then, the set $K=Q_{i}\setminus F$ intersects every quorum of $p_{j}$ and is therefore, a kernel of correct processes for it. Process $p_{j}$ will receive $\textsc{Readyafterecho},{n}]$m messages from $K$ and proceed to send a $\textsc{Readyafterready},{n,m}]$ messages.
Since all processes with depth 1 send a $\textsc{Readyafterready},{n,m}]$ message, all processes with depth 2 will receive such messages from one of their quorums, prompting them to send a
$\textsc{Readyafterecho},{n+1,m}]$ message. Since all processes with depth 2 will send such messages all processes with depth 3 will receive $\textsc{Readyafterecho},{n+1,m}]$ from one of their quorums and proceed to dar-deliver the value $m$.
Corollary 6.1.
Algorithm 1 solves RB[$d$] for any $d\geq 3$
Lemma 6.4.
In any execution with at least one process $p_{3}$ with depth 3 and for any $n>1$, if a process $p_{i}$ with depth 1 sends a $\textsc{Readyafterecho},{n,m}]$ message then there exists another process $p_{j}$ with depth 1 that sent a
$\textsc{Readyafterecho},{n-1,m}]$ message.
Proof.
According to the code of Algorithm 1, $p_{i}$ sends a
$\textsc{Readyafterecho},{n,m}]$ message after receiving
$\textsc{Readyafterready},{n-1,m}]$ messages from one of its quorums $Q_{i}\in\mathcal{Q}_{i}$. Process $p_{3}$ has a quorum $Q_{2}\in\mathcal{Q}_{3}$ such that all its members have depth at least 2. By the quorum consistency property there is at least one process $p_{2}\in Q_{i}\cap Q_{2}$. Therefore, $p_{2}$ sent a $\textsc{Readyafterready},{n-1,m}]$ message.
According to Algorithm 1, $p_{2}$ does this after receiving
$\textsc{Readyafterecho},{n-1,m}]$ messages from a kernel $K_{2}\in\mathcal{K}_{2}$. Since $p_{2}$ has depth 2, it has at least one quorum where all its members have depth 1, and therefore each of its kernels contains at least one process with depth 1. Let $p_{j}$ be such a process in $K_{2}$. Since $p_{j}$ has depth 1 and it sent a $\textsc{Readyafterecho},{n-1,m}]$ message we have proven the lemma. ∎
Lemma 6.5.
If a process $p_{i}$ with depth 1 sends a $\textsc{Readyafterecho},{1,m}]$ message then $\exists Q_{i}\in\mathcal{Q}_{i}$ such that all processes in $Q_{i}$ sent a $\textsc{Echo},{m}]$ message.
Proof.
According to Algorithm 1, a correct process only sends a
$\textsc{Readyafterecho},{1,m}]$ message after receiving $\textsc{Echo},{m}]$ messages from one of its quorums $Q_{i}\in\mathcal{Q}_{i}$. This proves the lemma. ∎
Lemma 6.6.
If a process $p_{i}$ with depth 3 delivers a value $m$ then there is a process $p_{j}$ with depth 1 such that $\exists Q_{j}\in\mathcal{Q}_{j}$ such that all processes in $Q_{j}$ sent a $\textsc{Echo},{m}]$ message.
Proof.
Let $p_{i}$ be a process with depth 3 that delivered a value $m$, and $Q_{2}\in\mathcal{Q}_{i}$ be the quorum of $p_{i}$ that contains only processes with depth at least 2. Its existence follows from the definition of depth. In order to deliver the value $p_{i}$ received $\textsc{Readyafterecho},{n,m}]$ messages from all processes in one of its quorums $Q_{i}\in\mathcal{Q}_{i}$ and some value $n$. Since $Q_{2}\cap Q_{i}\neq\emptyset$ there is at least one process $p_{j}^{\prime\prime}$ with depth 2 that sent a $\textsc{Readyafterecho},{n,m}]$ message.
Since a process with depth 2 also has depth 1, we can apply Lemma 6.4 to obtain that there is another process $p_{j}^{\prime}$ with depth 1 that sent a
$\textsc{Readyafterecho},{n-1,m}]$ message. By repeatedly applying Lemma 6.4 we reach a a process $p_{j}$ with depth 1 that sent a $\textsc{Readyafterecho},{1,m}]$ message. We now apply Lemma 6.5 to obtain that $\exists Q_{j}\in\mathcal{Q}_{j}$ such that all processes in $Q_{j}$ sent a $\textsc{Echo},{m}]$ message. This proves the lemma. ∎
In Algorithm 1 and Definition 6.1, we assume that all correct processes start the execution. This allows us to focus solely on characterizing the depth of the processes that complete the execution. However, in more complex scenarios, such as when protocols are composed sequentially, it is important to characterize the depth of the processes that start the execution.
For example, consider two protocols executed one after the other. If only processes with depth $d$ complete the first protocol, we cannot assume that all correct processes will initiate the second; we can only rely on those with depth $d$ to do so. This highlights the need to explicitly define both the starting and ending depths of participating processes, especially in composed protocols.
In the case of reliable broadcast, we treat it as a standalone algorithm and thus omit this distinction. However, in later sections, we will define protocols using two depth parameters: one for the processes that begin execution, and one for those that finish.
7 Consensus
We adapt the consensus problem to the depth paradigm and propose an algorithm that solves it for all processes with a minimum depth of at least 9. We follow an approach similar to the one proposed by Abraham et al. [1], using a round-based algorithm composed of a binding crusader agreement and a common coin. For this purpose, we introduce depth-characterized versions of these primitives in Sections 7.1 and 7.2, which we use as subprotocols inside the main consensus algorithm.
It is at this point that the need to characterize the depth of starting and finishing processes arises. The consensus protocol first calls the binding crusader agreement, and then the common coin. Only processes that finish the first call will be able to make the call to the second protocol. Because of this, we need to take into account that not all correct processes can start the execution. We introduce two parameters to characterize the depth of protocols: $d^{\prime}$ to refer to the depth of processes that start the execution, and $d$ to refer to the depth of processes that finish the execution. We will say that an algorithm works if, in case all processes with depth $d^{\prime}$ start the execution, all processes with depth $d$ satisfy the required properties. In case we assume that all correct processes start (i.e., $d^{\prime}=0$), as is the case for standalone protocols (like consensus and reliable broadcast), we will ignore the $d^{\prime}$ parameter.
7.1 Common Coin
The existing asymmetric common coin protocol proposed by Alpos et al. [2] has issues that motivate the development of a new protocol. It is not a purely asymmetric protocol, as each process first constructs the tolerated system, which is a symmetric fail-prone system, and executes the protocol based upon it. This means that the protocol is actually symmetric and is not suitable for asymmetric applications. We address this by first redefining the problem to fit the depth paradigm, shown in Definition 7.1, and propose a modified version of the scheme introduced by Alpos et al. [2] that solves the problem for all processes with depth at least 7, given that all processes with depth 6 start the execution. The parameter for all processes starting the protocol is needed since it will be used as part of a pipeline of protocols, which means that only the processes that finish the preceding protocol are guaranteed to start executing the common coin.
Definition 7.1 (Depth-characterized asymmetric common coin).
A protocol for common coin, for processes with depth $d^{\prime}$ and $d$, defined by the events release-coin(round) and output-coin(c, round), shortened as CC[d,d][$^{\prime},d$], satisfies the following properties:
Termination:
If all processes with depth $d^{\prime}$ call release-coin, then every process with depth $d$ eventually executes output-coin.
Unpredictability:
No process has information about the value of the coin before at least one process with depth $d^{\prime}$ has called release-coin.
Matching:
If a process with depth $d$ executes output-coin(c, round) and another process with depth $d$ executes output-coin(c, round), then $\textit{c}=\textit{c}^{\prime}$.
No bias:
The distribution of the coin is uniform over $\mathcal{B}$.
Algorithm 2 provides a solution to the common coin taking $d=7$ and $d^{\prime}=6$. Note that these values are not the smallest depths that allow the algorithm to solve the problem, but are the depths we will need within the context of the consensus algorithm. In fact, the same algorithm allows to solve CC[0, 1]. The common coin scheme follows the approach of Rabin [23] and assumes that coins are predistributed by a trusted dealer. The scheme uses Benaloh-Leichter [6] secret sharing, such that the coin is additively shared within every quorum. The dealer shares one coin for every possible round of the protocol and when a process receives a share it can verify that it was indeed created by the dealer. This can be done using standard digital signatures but is outside the scope of this paper. Alpos et al. [2] follow the same outline in their scheme.
State
$\mathbb{Q}$: asymmetric quorum system
$\textit{share}[Q][j]$: if $p_{i}\in Q$, this holds the share received from $p_{j}$ for quorum $Q$; initially $\bot$
upon event release-coin() do
for all $Q\in\mathcal{Q}_{j}$ where $\mathcal{Q}_{j}\in\mathbb{Q}$ such that $p_{i}\in Q$ do
let $s_{iQ}$ be the share of $p_{i}$ for quorum $Q$
for all $p_{j}\in\mathcal{P}$ do
send message $\textsc{share},{s_{iQ},Q}]$ to $p_{j}$
upon receiving a message $\textsc{share},{s,Q}]$ from $p_{j}$ such that $p_{j}\in Q$ do
if $\textit{share}[Q][j]=\bot$ then
$\textit{share}[Q][j]\leftarrow s$
upon exists $Q\in\mathcal{Q}_{i}$ such that for all $j$ with $p_{j}\in Q$, it holds $\textit{share}[Q][j]\neq\bot$ do
$s\leftarrow\sum_{j:p_{j}\in Q}\textit{share}[Q][j]$
output output-coin(s)
Algorithm 2 Asymmetric common coin (code for $p_{i}$)
Lemma 7.1.
Consider an execution with faulty processes $F$, a process $p_{i}$ such that $\textsl{depth(}p_{i}\textsl{)}\geq 1$ and one of its quorums $Q_{i}\in\mathcal{Q}_{i}$. For any other process $p_{j}$ with $\textsl{depth(}p_{j}\textsl{)}\geq 1,\exists K_{j}\in\mathcal{K}_{j}$ such that $K_{j}\subseteq Q_{i}\setminus F$.
Proof.
Since $p_{i}$ and $p_{j}$ both have depth at least 1, it holds $F\in\mathcal{F}_{i}^{*}$ and $F\in\mathcal{F}_{j}^{*}$. This implies $F\in\mathcal{F}_{i}^{*}\cap\mathcal{F}_{j}^{*}$. Then, the set $Q_{i}\setminus F$ intersects every quorum of $p_{j}$ by the quorum consistency property, and therefore contains a kernel for $p_{j}$. ∎
Lemma 7.2.
Algorithm 2 implements a depth-characterized common coin for $d=7$ and $d^{\prime}=6$.
Proof.
Termination: Consider a process $p_{i}$ with depth 7. There exists a quorum $Q\in\mathcal{Q}_{i}$ such that all its members have depth at least 6. Since we assume that they all start the protocol, they will release the coin and send their share to all the processes in $\mathcal{P}$. Therefore, $p_{i}$ will receive all the coin shares for at least the quorum $Q$ and will be able to output the coin value.
Unpredictability: Assume a process $p_{7}$ with depth 7 outputs the coin. This implies the existence of a set $Q\in\mathcal{Q}_{7}$ such that each member of $Q$ sent a share message. Since $p_{7}$ has a quorum $Q_{7}\in\mathcal{Q}_{7}$ such that $\forall p\in Q_{7},\textsl{depth(}p\textsl{)}\geq 6$ and $Q\cap Q_{7}\neq\emptyset$, it follows that at least one process with depth 6 has released the coin.
Matching: Follows from the fact that the value of the coin for each round is predetermined by the dealer.
No bias: Follows from the fact that the value of the coin for each round is predetermined by the dealer. ∎
7.2 Binding Crusader Agreememt
Following the paradigm of Abraham et al. [1], we will build the consensus protocol relying on binding crusader agreement primitive. Definition 7.2 presents the properties that such a primitive must satisfy. We adapt the properties to reflect that only a fraction of correct processes might start the execution. This is of use in the round-based context where this primitive will be used, as we can only guarantee that processes with a certain depth will be alive in consensus rounds other than the first.
Definition 7.2 (Depth-characterized asymmetric binding crusader agreement).
A protocol for binding crusader agreement with depths $d^{\prime}$ and $d$, defined through the events bca-propose(v, round) and bca-decide(v, round), shortened as BCA[d,d][$^{\prime},d$], satisfies the following properties:
Termination:
If all processes with depth $d^{\prime}$ bca-propose a value then every process with depth $d$ eventually bca-decides.
Validity:
If all processes with depth $d^{\prime}$ bca-propose the same input value $v$, then this is the only possible bca-decide output for all processes with depth $d$.
Agreement:
If two processes with depth $d$ bca-decide $v$ and $v^{\prime}$, then $v=v^{\prime}$ or one of the values is $\bot$.
Binding:
Let time $\tau$ be the first time at which there is a process with depth $d$ that bca-decides at time $\tau$. At time $\tau$, there is a value $b\in\{0,1\}$ such that no party with depth $d$ bca-decides 1b1-b in any extension of this execution.
Algorithm 3 presents a solution to this problem for $d^{\prime}=2$ and $d=6$. We prove its correctness in Lemma 7.6. Lemmas 7.37.4, and 7.5 prove helper results which are of use in the proof of the main lemma.
State
$\textit{sentecho2}\leftarrow\textsc{false}$: variable to control the sending of messages
$\textit{approved}\leftarrow\{\}$: array to store potential output values
upon event bca-propose(input) do
send message $\textsc{echo},{\textit{input}}]$ to all
upon receiving a message $\textsc{echo},{\textit{c},\textit{v}}]$ from all processes in some $K\in\mathcal{K}_{i}$ do
send message $\textsc{echo${}^{\prime}$},{\textit{c},\textit{v}}]$ to all
upon receiving a message $\textsc{echo${}^{\prime}$},{\textit{c},\textit{v}}]$ from all processes in some $Q\in\mathcal{Q}_{i}$ do
send message $\textsc{echo},{\textit{c}+1,\textit{v}}]$ to all
upon receiving a message $\textsc{echo},{\textit{c},\textit{v}}]$ from all processes in some $Q\in\mathcal{Q}_{i}$ do
$\textit{approved}\leftarrow\textit{approved}\cup\{v\}$
if $\neg\textit{sentecho2}$ then
send message $\textsc{echo2},{\textit{v}}]$ to all
$\textit{sentecho2}\leftarrow\textsc{true}$
upon receiving a message $\textsc{echo2},{\textit{v}}]$ from all processes in some $Q\in\mathcal{Q}_{i}$ or $|\textit{approved}|>1$ do
if $|\textit{approved}|>1$ then
send message $\textsc{echo3},{\bot}]$ to all
else
send message $\textsc{echo3},{\textit{v}}]$ to all
upon receiving a message $\textsc{echo3},{\textit{v}}]$ from a $Q\in\mathcal{Q}_{i}$ or $(|\textit{approved}|>1\text{ and received {echo3} messages from a }Q\in\mathcal{Q}_{i})$ do
if received $\textsc{echo3},{v}]$ from a $Q\in\mathcal{Q}_{i}$ then
bca-decide($v$)
else
bca-decide($\bot$)
Algorithm 3 Asymmetric binding crusader agreement (code for $p_{i}$)
Lemma 7.3.
Given an execution with faulty processes $F$, a process $p_{i}$ with depth 2, one of its quorums $Q\in\mathcal{Q}_{i}$, and any possible mapping $M:\mathcal{P}\rightarrow\{0,1\}$. There is a value $v\in\{0,1\}$ such that for any other process $p_{j}$ with depth 1, $\exists K\in\mathcal{K}_{j}$ such that $K\subseteq Q\setminus F$ and $\forall p\in K,M(p)=v$
Proof.
We will prove the lemma with the help of a simple algorithm. Each process $p_{i}$ starts the execution with an initial value $v_{i}\in\{0,1\}$, which is sent to all other processes. Afterwards, each process collects values coming from the processes in one of their quorums $Q_{i}\in\mathcal{Q}_{i}$ in a set denoted $V_{i}$.
Suppose that all correct processes execute this algorithm and consider any possible assignment of initial values to the processes, the mapping $M$. We will show that for any process with depth 1 $p_{i}$, there is a value $v$ and a set of correct processes $K\subseteq Q_{i}$ such that all processes in $K$ sent a message with value $v$ and $K$ is a kernel for all processes with depth 1.
$p_{i}$ received messages from $Q_{i}$, let $F$ be the faulty parties in the execution and let $S_{v^{\prime}}\subseteq Q_{i}$ denote the processes in Q which sent value $v^{\prime}$. Suppose $Q_{i}\setminus F$ has no kernel for all other processes with depth 1 such that all of them send the same value $v$. This means that there exists a process with depth 2 $p_{1}$ and a quorum $Q_{1}\in\mathcal{Q}_{1}$ such that $S_{0}\cap Q_{1}=\emptyset$, that is, all correct processes in $Q_{1}$ sent the value 1. This means that the processes in $S_{0}$ cannot form a kernel for all other processes. Similarly, assume that there is a process with depth 1 $p_{0}$ and a quorum $Q_{0}\in\mathcal{Q}_{0}$ such that $S_{1}\cap Q_{0}=\emptyset$. However, if this occurs, $Q_{0}\cap Q_{1}\subseteq F$ which contradicts the consistency property of the quorum system since $p_{0}$ and $p_{1}$ both have depth 1.
This holds for all quorums of all processes with depth 1 and any mapping $M$. Therefore, any quorum contains a kernel with the same value for all other processes with depth 1. This will be of use when proving the termination property of Algorithm 3. ∎
Lemma 7.4.
If all processes with depth 2 start Algorithm 3 with the same input value $v$, all processes with depth 6 will output $v$.
Proof.
The algorithms is used to solve C[9], so we will assume that the execution contains at least one process $p_{8}$ with depth 8. This process has a quorum $Q_{8}$ such that all its members have depth at least 7. Suppose that a process $p_{4}$ with depth 4 receives echo messages with value 1v1-v from one of its quorums $Q_{4}\in\mathcal{Q}_{4}$. Since $Q_{4}\cap Q_{8}\neq\emptyset$, there is at least one process $p_{7}$ with depth 7 that sent an echo message with value 1v1-v towards $p_{4}$. However, $p_{7}$ started with value $v$, so it must have received echo messages with value 1v1-v from one of its quorums $Q_{6}\in\mathcal{Q}_{6}$ (line 3). Since $Q_{6}\cap Q_{7}\neq\emptyset$, there is a process $p_{6}^{\prime}$ with depth 6 that received echo messages with value 1v1-v from one of its kernels $K_{6}^{\prime}\in\mathcal{K}_{6}^{\prime}$ (line 3). Since any kernel for $p_{6}^{\prime}$ must contain at least one process with depth 5, there is a process $p_{5}$ with depth 5 that sent an echo message with value 1v1-v. Using the same logic as before, $p_{5}$ must have received a quorum of echo messages from some $Q_{5}\in\mathcal{Q}_{5}$. Since $Q_{5}\cap Q_{7}\neq\emptyset$, some process $p_{6}^{\prime\prime}$ with depth 6 must have received echo messages with value 1v1-v from one of its kernels. Following this reasoning, we reach an infinite sequence of processes with depth 6 that must have received echo messages with value 1v1-v from one of their kernels. But since no process with depth 2 has initial value 1v1-v, we can conclude that this is not possible and therefore no process with depth 4 receives a quorum of echo messages with value 1v1-v.
We have established that all processes with depth 4 will only receive a quorum of echo messages with value $v$. Therefore, they will send a $\textsc{echo2},{v}]$ message (line 3 and all processes with depth 5 will receive a quorum of this messages, causing them to send a $\textsc{echo3},{v}]$ message. Finally, all processes with depth 6 will receive a quorum of these messages and deliver value $v$ (line 3).
Lemma 7.5.
If a process with depth 5 sends a $\textsc{echo3},{v}]$ message in Algorithm 3, no other process with depth 5 will send a $\textsc{echo3},{1-v}]$ message.
Proof.
Assume that process $p_{5}$ with depth 5 is the first process with depth 5 to send a echo3 message and does so with value $v$. To do this, it must have received echo2 messages with value $v$ from one of its quorums $Q_{5}\in\mathcal{Q}_{5}$(line 3). Imagine that there is another process $p_{5}^{\prime}$ with depth 5 that sends a $\textsc{echo3},{1-v}]$ message. This requires receiving $\textsc{echo2},{1-v}]$ messages from one of its quorums $Q_{5}^{\prime}\in\mathcal{Q}_{5}^{\prime}$. Since $(Q_{5}\cap Q_{5}^{\prime})\setminus F\neq\emptyset$, and a nonfaulty party only sends echo2 messages with one value, $p_{5}^{\prime}$ could not have sent a $\textsc{echo2},{1-v}]$ message. ∎
Lemma 7.6.
Algorithm 3 solves asymmetric binding crusader agreement with $d=6$ and $d^{\prime}=2$.
Proof.
Termination:
We assume that all processes with depth $d^{\prime}=2$ start the execution. Upon starting they send a a $\textsc{echo},{1,v}]$ message with their respective initial value $v$ (line 3). Since all processes with depth $d^{\prime}+1$ have a quorum formed by processes of depth $d^{\prime}$, they will eventually receive $\textsc{echo},{1,*}]$ messages from one of their quorums. Some of these messages will have $v=0$ and others $v=1$. By Lemma 7.3 we know that there is a value $v\in\{0,1\}$ such that all processes in the quorum which sent $v$ are also a kernel for all other processes with depth 2. Therefore, all processes with depth 2 will receive echo messages with the same value from one of their kernels.
This causes them to send a $\textsc{echo${}^{\prime}$},{1,v}]$ message (line 3). Since all processes with depth 2 will send this message, all processes with depth 3 are guaranteed to receive $\textsc{echo${}^{\prime}$},{1,v}]$ from one of their quorums. This triggers them to send a $\textsc{echo},{2,v}]$ message (line 3). Since all depth $d^{\prime}+1=3$ processes send this message, all depth $d^{\prime}+2$ messages will receive a quorum of $\textsc{echo},{2,v}]$ messages, triggering the sending of a $\textsc{echo2},{v}]$ message if none has been sent so far (line 3). Consequently, all processes with depth $d^{\prime}+3$ will receive a quorum of $\textsc{echo2},{v}]$ messages, send a $\textsc{echo3},{v}]$ message, and all processes with depth $d^{\prime}+4$ will receive a quorum of $\textsc{echo3},{v}]$ messages. This will trigger them to deliver a value, $v$ or $\bot$, depending on the messages received. This shows that all processes with depth 6 will deliver a value.
Agreement: Suppose processes $p_{0}$ and $p_{1}$, both with depth $d$, output values 0 and 1 respectively. This means that $p_{0}$ received $\textsc{echo3},{0}]$ messages from one of its quorums $Q_{0}\in\mathcal{Q}_{0}$ and $p_{1}$ received $\textsc{echo3},{1}]$ messages from one of its quorums $Q_{1}\in\mathcal{Q}_{1}$. If $d\geq 1$, quorum consistency guarantees that $Q_{0}\cap Q_{1}\nsubseteq F$. Let $p_{k}$ be one correct process in $Q_{0}\cap Q_{1}$. Since $p_{0}$ and $p_{1}$ delivered different values, $p_{k}$ must have sent $\textsc{echo3},{0}]$ to $p_{0}$ and $\textsc{echo3},{1}]$ to $p_{1}$. But since $p_{k}$ is correct and the algorithm states that each process sends only one echo3 message this is not possible.
Validity: Follows from Lemma 7.4.
Binding: Consider the first process $p_{6}$ with depth $d=6$ that sees Echo3 messages from one of its quorums $Q_{6}\in\mathcal{Q}_{6}$. There are two cases, if a process with depth 5 sent a non-$\bot$ message then we have binding from Lemma 7.5. Otherwise, all processes in $Q_{6}\setminus F$ sent a $\textsc{echo3},{\bot}]$ message. From Lemma 7.1 we know that $Q_{6}\setminus F$ has a kernel for any other process $p_{6}^{\prime}$ with depth 6. Therefore, it is not possible for $p_{6}^{\prime}$ to see a quorum of Echo3 messages for any non-$\bot$ value.
We have shown that, if we take $d^{\prime}=2$ and $d=6$, the algorithm satisfies all required properties. Therefore, Algorithm 3 solves asymmetric binding crusader agreement with $d=6$ and $d^{\prime}=2$. ∎
7.3 Consensus
Definition 7.3 introduces the concept of depth to the asymmetric consensus problem. We propose a consensus algorithm using the round-based paradigm of Abraham et al. [1]. Each round consists of a call to the binding crusader agreement and common coin primitives presented earlier. If the values obtained from the two primitives match, then the processes deliver it; otherwise, another round is executed.
We make a small addition to the common coin and binding crusader agreement calls. When one of these sub-protocols is triggered, it is done so with a round tag. This is used to keep track of the round that is running and does not interfere with any details of the sub-protocols.
Definition 7.3 (Depth-characterized asymmetric binary Byzantine consensus).
A protocol for Byzantine consensus, shortened as C[d], defined through the events c-propose(val) and c-decide(val), satisfies:
Probabilistic termination:
Every process with depth $d$ eventually c-decides with probability $1$.
Validity:
If all correct processes c-propose the same value val, then all processes with depth $d$ must c-decide val.
Agreement:
No two processes with depth $d$ c-decide different values.
A difference with respect to symmetric round-based protocols is that not all processes will finish all the sub-protocol calls. Only processes with depth 6 can complete the crusader agreement, and only those with depth 7 also finish the subsequent common coin. If we take a naive approach, executing rounds one after another, where only those processes that completed the previous round participate in the next, the required depth increases with the number of rounds. Depth 7 is required to finish the first round, 14 for the second round, 21 for the third and so on.
To fix this, we introduce a revival mechanism, which lets any process with depth at least 2 rejoin future rounds if it ever gets stuck. Without this element, after each round, liveness could only be guaranteed for processes with higher and higher depths. Since such a protocol can have executions with infinite length, we would only be able to provide guarantees for processes with infinite depth, which goes against our objectives here.
The revival mechanism is similar to the quorum amplification technique proposed by Bracha [7]. Each process that finishes the pipeline broadcasts a revive1 message. If termination for a round is guaranteed for processes with depth $7$ but there is also a process with depth $9$, Lemma 7.3 implies that there will be a kernel for all processes with depth 1 formed by processes that finish the pipeline with the same value. Upon receiving $\textsc{revive1},{v,round}]$ messages from a kernel, each process sends a $\textsc{revive2},{v,round}]$ message. We know that all processes with depth at least 1 will send this message. Afterwards, upon receiving revive2 messages from all processes in a quorum, the process will go back to the beginning of the round pipeline. This will happen for all processes with depth 2. Therefore, if all processes with depth $7$ finish the pipeline and we also have at least one process with depth $9$, we can revive all processes with depth 2. This is the reason why our protocol solves C[9] but not C[7], at least one process with depth 9 is needed to make the revival mechanism work.
A process may execute multiple rounds in parallel. If the revival mechanism signals a process to start the next round, while the current round is still ongoing, the process proceeds with both executions simultaneously. This is enabled by the usage of arrays (v, continue) to store the process states separately for each round. This design ensures that the process can continue sending messages for every round it has initiated, which may be needed by other processes in the system.
Algorithm 4 presents a solution to the asymmetric consensus problem for all processes with depth 9. It relies on the binding crusader agreement and common coins previously introduced. Lemma 7.13 proves that it satisfies the properties described in Definition 7.3. Since we want to prove the properties for all processes with depth 9, in the proofs, we will assume that there is at least one such process.
Lemma 7.7.
If a process with depth at least 7 has the variable continue[round+1] set to true (i.e. start the round execution) in Algorithm 4, then at least one processes with depth at least 7 has continue[round] set to true (i.e., finishes the previous round execution).
Proof.
We are solving C[9], therefore, we will assume the existence of at least one process $p_{9}$ with depth 9.
Assume that process $p_{7}$ sets continue[round+1] to true. There are two ways in which this can happen. It can either receive a quorum of $\textsc{revive2},{val,round+1}]$ messages (line 4) or obtain an output from the common coin (line 4. In the second case, it is clear that at least one process with depth 7 has $\textit{continue}[\textit{round}]=\textsc{true}$, as $p_{7}$ itself satisfies this. This follows from the fact that for $p_{7}$ to set the variable to true, it must have obtained an output from the common coin (line 4) and before that from the binding crusader agreement (line 4). To obtain a bca-decide, it must have first called bca-propose, which only happens if v[continue] is true.
In the other case, if it received $\textsc{revive2},{val,round+1}]$ messages from a quorum $Q_{7}\in\mathcal{Q}_{7}$, we know that all members of $Q_{7}$ received a kernel of $\textsc{revive1},{val,round+1}]$ messages. Let $Q_{9}\in\mathcal{Q}_{9}$ be the quorum of $p_{9}$ such that all its members have depth at least 8 and denote by $p_{8}$ a process that belongs to $Q_{7}\cap Q_{9}$. Since $p_{8}$ has depth 8 and it received revive1 messages from one of its kernels $K_{8}\in\mathcal{K}_{8}$, we know that at least one process $p_{7}^{\prime}$ with depth 7 sent a $\textsc{revive1},{val,round+1}]$ message. By the argument used in the previous paragraph, we know that a process only sends this message if it has $\textit{v}[\textit{round}]=\textsc{true}$. This proves the lemma. ∎
Lemma 7.8.
If all processes with depth at least 2 have the variable continue[round] set to true (i.e. start the round execution) in Algorithm 4, then all processes with depth at least 7 will set continue[round+1] to true (i.e., finish the round execution).
Proof.
A process with depth at least 7 sets [round+1] to true upon finishing the execution of a round (line 4). From the termination properties of the binding crusader agreement and the common coin we know that if all processes with depth 2 start a round, then all processes with depth 6 will finish the binding crusader agreement (i.e., reach line 4) and all processes with depth 7 will finish the common coin (i.e., finish line 4). All correct processes that finish the common coin call will also finish the round, as there is no other point in which they can be stopped between lines 4 and 4. Therefore, all processes with depth 7 will set continue[round+1] to true and continue executing the next round. This also implies that all processes with depth at least 7 will send a $\textsc{revive1},{v,round}]$ message for some value v. ∎
Lemma 7.9.
If all processes with depth at least 7 set continue[round+1] to true in line 4 of Algorithm 4, then all processes with depth 2 will eventually have $\textit{continue}[\textit{round}+1]$ also set to true.
Proof.
We are solving C[9], therefore, we will assume the existence of at least one process $p_{9}$ with depth 9. Since all processes with depth at least 7 send a revive1 message for round $\textit{round}+1$, $p_{9}$ will receive such messages from all processes in one of its quorums $Q_{9}\in\mathcal{Q}_{9}$. By Lemma 7.3, we know that there will be a value val and a subset of correct processes $K\subseteq Q_{9}$ such that all processes in $K$ sent $\textsc{revive1},{val,round+1}]$ and $K$ is a kernel for all processes with depth 1. Therefore, all processes with depth 1 will send $\textsc{revive2},{val,round+1}]$ message (line 4) and all processes with depth 2 will receive a quorum of such messages. This triggers them to enter round $r+1$ by setting the variable $\textit{continue}[\textit{round}+1]$ to true and the variable $\textit{v}[\textit{round}+1]$ to the value from the kernel. ∎
Lemma 7.10.
If all processes with depth at least 7 send a $\textsc{revive1},{val,round}]$ message with the same value val, then all processes with depth at least 2 will receive a $\textsc{revive2},{val,round}]$ message from one of their quorums and will not receive $\textsc{revive2},{1-val,round}]$ messages from one of their quorums.
Proof.
Since we are providing a solution for C[9] we assume that there exists at least one process $p_{9}$ with depth 9. Since all processes with depth 7 send this message, $p_{9}$ will receive $\textsc{revive1},{val,round}]$ from at least one of its quorums $Q_{9}\in\mathcal{Q}_{9}$. From Lemma 7.1, we know that $Q_{9}\setminus F$ is a kernel for all processes with depth at least 1, causing all such processes to send a $\textsc{revive2},{val,round}]$ message (line 4). All processes with depth at least 2 will receive this message from one of their quorums.
We now show that no process with depth at least 2 will receive $\textsc{revive2},{1-val,round}]$ messages from all processes in one of its quorums. Suppose that one such process $p_{2}$ receives these messages from a quorum $Q_{2}\in\mathcal{Q}_{2}$. Since $Q_{2}\cap Q_{9}\neq\emptyset$, there is at least one process $p_{8}$ with depth 8 that sent a $\textsc{revive2},{1-val,round}]$ message. $p_{8}$ would have done this after receiving $\textsc{revive1},{1-val,round}]$ from one of its kernels $K_{8}\in\mathcal{K}_{8}$. From the definition of depth we know that $p_{8}$ has a quorum $Q_{8}\in\mathcal{Q}_{8}$ such that all its members have depth at least 7. Since $K_{8}\cap Q_{8}\neq\emptyset$, this means that a process with depth 7 must have sent a $\textsc{revive1},{1-val,round}]$ message, but by the lemma assumptions this is not possible.
Since all processes with depth at least 2 receive a quorum of $\textsc{revive2},{val,round}]$ messages, but no quorum of $\textsc{revive2},{1-val,round}]$ messages, we have proven the lemma. This also implies that they will set v[round] to the value val and continueround] to true (line 4), enabling them to continue execution in round round. ∎
Lemma 7.11.
If a process with depth at least 2 receives $\textsc{revive2},{v,round}]$ messages from one of its quorums, then at least one process with depth at least 7 sent a $\textsc{revive1},{v,round}]$ message.
Proof.
As we did before, we assume that there exists at least one process $p_{9}$ with depth 9. Suppose a process $p_{2}$ with depth 2 received $\textsc{revive2},{v,round}]$ messages from one of its quorums $Q_{2}\in\mathcal{Q}_{2}$. Since $Q_{2}\cap Q_{9}\neq\emptyset$, this means that at least one process $p_{8}$ with depth 8 sent a $\textsc{revive2},{v,round}]$ message. Following the algorithm, this only happens if $p_{8}$ received $\textsc{revive1},{v,round}]$ messages from one of its kernels $K_{8}\in\mathcal{K}_{8}$ (line 4). Since $p_{8}$ has at least one quorum composed only of processes with depth 7, $K_{8}$ contains one such process. Therefore, a process with depth 7 sent a $\textsc{revive1},{v,round}]$ message. ∎
Lemma 7.12.
If all processes with depth at least 2 have $\textit{v}[\textit{round}]$ set to the same value val in Algorithm 4, then all processes with depth 9 will c-decide val within a constant expected number of rounds.
Proof.
From Lemma 7.4 we know that all processes with depth 6 will obtain the same value val as a result of the binding crusader agreement. We also know that all processes with depth at least 7 will output a common coin value c. If the random value of the coin c matches val, then all processes with depth 9 will decide val in this round. Otherwise, by Lemma 7.10, all processes with depth 2 will start the next round with value val. This continues until a round is reached in which $\textit{c}=\textit{val}$, which triggers all processes with depth at least 9 to c-decide val. This happens with probability frac12frac{1}{2} in each round. This follows a geometric distribution and will therefore occur within an expected constant number of rounds. ∎
Lemma 7.13.
Algorithm 4 solves C[9].
Proof.
Probabilistic Termination: If in a round $r$ all processes with depth 6 output the same value, $\bot$ or val, in their binding crusader agreement, by Lemma 7.12, the execution will finish in an expected constant number of rounds.
Consider the first round of the execution. if all processes with depth at least 6 output the same value, $\bot$ or val, in their binding crusader agreement, by Lemma 7.12 the execution will finish in an expected constant number of rounds.
Otherwise, by the binding property of the crusader agreement, the adversary is bound to some non-$\bot$ value $v$ by the time the first process with depth 6 finishes the crusader agreement. By the unpredictability property of the common coin, this must happen before the coin value c is revealed. Therefore, with probability $\frac{1}{2}$, $\textit{val}=\textit{c}$ and all processes with depth at least 7 set v[round+1] to value val. Then, by applying Lemma 7.10 and then Lemma 7.12, we conclude that within a constant expected number of rounds, a round will be reached where all processes with depth 9 will decide. Otherwise, some processes with depth 7 will set v[round+1] to value val and others will set it to the value c of the coin. By Lemma 7.8 we know that all of them will set continue[round+1] to true. Applying Lemma 7.9 we know that eventually, all processes with depth at least 2 will set continue[round+1] to true and continue the execution in the next round.
The next round presents the same situation, where some processes with depth at least 2 start round+1 with value 0 and others do it with value 1. This will occur until the value of the coin matches the output of the binding crusader agreement, which will also take a constant expected number of rounds.
Agreement: Consider the first round $r$ in which a process decides a value $b$. The value $b$ must be the output from both the binding crusader agreement and the common coin (line 4). Since the value decided must match the coin, any process that decides in round $r$ will also decide $b$. All processes that complete this round will do so with the value $b$ stored in the variable v. Processes that obtain $\bot$ from the crusader agreement will set v to be the value of the coin (line 4), while all other processes will set v to be the output of the crusader agreement (line 4). Therefore, all processes with depth 7 finish the round with $b$ stored in the variable $v$, and when processes with depth 2 are revived, following Lemma 7.10, they do so with value $b$ in their variable v.
We know from Lemma 7.11 that a process with depth 2 will not start round round+1 until at least one process with depth 7 has finished the previous round. Therefore, all processes with depth 2 start the next round and do so with value $b$. Lemma 7.4 shows that all processes with depth 6 will output $b$ from the crusader agreement, and as such no process will decide 1b1-b in the round. If a decision takes place in this round it will be with value $b$, otherwise, all processes with depth 2 start the next round with value $b$. This continues for all subsequent rounds and as a consequence agreement holds.
Validity: We assume that all correct processes (i.e., those with depth 0) start the execution with the same value. Therefore they will set $\textit{v}[1]$ to val. By Lemma 7.12 we know that all processes with depth 9 will c-decide val.
State
$\textit{v}[\ ]$: array that stores input values for each round, initially empty
$\textit{continue}[\ ]\leftarrow[\textsc{false},\dots,\textsc{false}]$: array for flow control, initially all false
upon event c-propose(value) do
v[1] $\leftarrow$ value
$\textit{continue}[1]\leftarrow\textsc{true}$
upon exists round such that $\textit{continue}[\textit{round}]$ do
continue[round] $\leftarrow\textsc{false}$
bca-propose(v[round]) with tag round// calls the BCA protocol
upon bca-decide(value) with tag round do // after finishing the BCA protocol
v[round+1] $\leftarrow$ value
release-coin() with tag round // call the common coin protocol
upon output-coin(c) with tag round do // after finishing the common coin protocol
[…]
if c = v[round+1] then
c-decide(v[round+1])
else
v[round+1] $\leftarrow$ c
send message $\textsc{revive1},{\textit{v}[\textit{round}+1],\textit{round}}]$ to all $p\in\mathcal{P}$
$\textit{continue}[\textit{round}+1]\leftarrow\textsc{true}$
upon receiving $\textsc{revive1},{\textit{v${}^{\prime}$},\textit{round}}]$ from all processes in some $K\in\mathcal{K}_{i}$ do
send message $\textsc{revive2},{\textit{v${}^{\prime}$},\textit{round}}]$ to all $p\in\mathcal{P}$
upon receiving $\textsc{revive2},{\textit{v${}^{\prime}$},\textit{round}}]$ from all processes in some $Q\in\mathcal{Q}_{i}$ do
$\textit{v}[\textit{round}]\leftarrow\textit{v}^{\prime}$
$\textit{continue}[\textit{round}]\leftarrow\textsc{true}$
Algorithm 4 Asymmetric binary consensus (code for $p_{i}$)
8 Conclusions and Future Work
Prior to this work, solutions to problems such as reliable broadcast and consensus typically relied on the existence of a guild. Theorem 4.1 shows that this assumption effectively reduces to the existence of a symmetric system, constructed from the original assumptions, that provides equivalent or stronger guarantees. In other words, existing solutions are not well-suited for genuinely asymmetric settings.
To overcome this limitation, we introduced a more fine-grained way to characterize asymmetry through the notion of depth. We proved that reliable broadcast can be solved for all processes with depth at least 3, and consensus for processes with depth at least 9. This significantly relaxes the requirements compared to prior work by Alpos et al. [2], which established feasibility only for processes with infinite depth.
The exact minimum depth required for solving asymmetric reliable broadcast and consensus remains an open problem. For reliable broadcast, the lower bound is either 2 or 3; for consensus, it lies between 1 and 9. Determining these thresholds is crucial, as they capture the precise necessary and sufficient conditions for solving problems under asymmetric trust. Previous work [2, 17] established sufficiency but left necessity unproven.
Finally, an important direction for future research is to explore whether the concepts and techniques introduced here extend to other asymmetric trust models [17, 19].
Acknowledgments
This work was supported by the Swiss National Science Foundation (SNSF) under grant agreement Nr. 219403 (Emerging Consensus), the Cryptographic Foundations for Digital Society, CryptoDigi, DFF Research Project 2, Grant ID 10.46540/3103-00077B, and the Initiative for Cryptocurrencies and Contracts (IC3).